***PLEASE SHOW ALL WORK AND COMMANDS FOR HOW TO DO IN SAGE APPLICATION***

The first nine digits of the ISBN-10 of the European version of the fifth edition of this book are 0-07-119881.


What is the check digit for that book?
Answer #1

Check digits for ISBN10 satisfy: Sum of products of digits times position (taken from right) is 0 modulo 11

So if the check digit is x then we must have

0 x 10+0 x 9+7x 8+1x7+1x6+9x5+8 x 4+8x3+1x2+1x1 = 0 (mod 11)

So that 2+172 = 0 (mod 11) ►r= -172 (mod 11) = 4

Required check-digit is {\color{Red} 4}

\blacksquare

Hope this was helpful. Please do leave a positive rating if you liked this answer. Thanks and have a good day!